I need to know how to do this question. It is in fraction form though
3(x+5)= 3
7(x-2)= 5

Answer by stanbon(75887)

3/(x+5)= 3
Cross-multiply to get:
3(x+5) = 3
x+5 = 1
x = -4
===================
7/(x-2)= 5
Cross-multiply to get:
5(x-2) = 7
5x-10 = 7
5x = 17
x = 3.4
============
Cheers,
Stan H.
============